OpenStudy (anonymous):

Look at this linear equation. 3x + 5y = 30 Which of the following equations can be used to solve for y in the linear equation above?

OpenStudy (anonymous):

\[ y = 6-3/5\cdot x \]

OpenStudy (anonymous):

y=(30-3x)/5

OpenStudy (anonymous):

A. y = -3/5 x + 30 B. y = 3x + 6 C. y = - 3/5 x + 6 D. y = 3/5 x + 30

OpenStudy (anonymous):

C

OpenStudy (sriram):

C

OpenStudy (anonymous):

C

OpenStudy (anonymous):

http://youthvisit.com/?userid=1578